Molander Corporation is a distributor of a sun umbrella used at resort hotels. Data concerning the next month’s budget appear below:

Selling price per unit $ 23

Variable expense per unit $ 14

Fixed expense per month $ 7,290

Unit sales per month $ 960

1. What is the company’s margin of safety? (Do not round intermediate calculations.)

2. What is the company’s margin of safety as a percentage of its sales? (Round your percentage answer to 2 decimal places (i.e. 0.1234 should be entered as 12.34).)
Answer #1

Contribution margin=Sales-Variable cost  

=(23-14)=$9 per unit

Breakeven=Fixed expenses/Contribution margin

=(7290/9)=810 units

Margin of safety=Total sales-Breakeven sales

=960-810

=150 units

=(150*23)

=$3450

Margin of safety%=Margin of safety/Total sales

=150/960

=15.63%(Approx)